# A Coalition Game for On-demand Multi-modal 3D Automated Delivery System

### v1

@@ 5.5.1 Coalition Generalizability
The overall result shows an acceptable solution generated from each of the scales to others; most notably, the generalizing on smaller networks from the trained model of larger networks produces better results. This is probably due to the inclusion of smaller and more information on the network embedding. This variation can also be explained by the different number of depots in each case, which causes the training parameter not to capture the depots embedding well. In addition, as the scale of the problem gets larger, generalization results undergo variability, particularly due to managing the fewer vehicles within the larger network and generalizing from lower vehicle-to-request ratio distribution, which causes ineffectiveness in high request densities assignment with limited resources. Nevertheless, in the constant size of the graph, the model’s ability to generalize is still effective, and in the next part, we analyzed the core game theory for the constant size of the network, yet various number of agents.
@@ 5.5.2 Coalitional Ability
We use the pre-trained model for E-CPDPTW120-d3 for ten agents (five of each mode) to conduct a test to find the core coalition among the subset of agents. To evaluate the cost function for each coalition, we use the generalizing on the pre-trained model to obtain cost values to avoid the high computation expense of training for each coalition. The coalitions are of the two separate modes of ADRs ($r\in\left\{1,\ldots,N^{r}\right\}$) and UAVs ($d\in\left\{1,\ldots,N^{d}\right\}$) set, assumed to have no intersection; thereby, we would then consider a convex game according to Shapley (1971). This refers to Equation 29, a property of convex game for coalitions formed of cooperation.
$C(S_{1}\cup S_{2})\leq C(S_{1})+C(S_{2})-C(S_{1}\cap S_{2})\;\;\forall S_{1},S_{2}\subseteq N$ (29)
@@
Since the formation is under two completely different coalitions, we conclude that $S_{1}\cap S_{2}=\emptyset$; therefore, the convex property of the game is met. The Algorithm 1 is used here to assess how coalitions will perform and if the core exists. For up to a group of just consisting of either UAVs or ADRs, the individual cost of both modes will calculated, as well as their coalition. Subsequently, the sub-additivity and efficient conditions are checked so that not only must the cooperative cost be less than both individual modes, but also, the newly allocated cost of the coalition must be less than the average mode cost alone. Figure 16 displays contour plots for coalition gains; the heat map bar shows the difference between the sum of individual and cooperative modes. Note that the spectrum level is set to be sufficiently high to obtain a smooth colour transition graph. However, it is clear that the core analysis plot is discrete. The star sign in the figure indicates if the core exists, given its criteria. It was observed in both scales of the problem that the core exists except in the network size of 120, where only one UAV is operating. This can be mainly because in the larger network, having one UAV independently from the ADR numbers cannot accommodate the demand due to high delays in the deliveries of ADRs, yet it is achievable in the $n=50$. Nonetheless, lower coalition values are observed in both graphs where fewer UAVs are involved, and the lowest value gain is obtained, which fails to perform well in such a large network. It is, however, toward yellowish regions where the most value is gained by acting cooperatively instead, which is increasing where both the number of UAVs and ADRs increase. This implies that a balanced configuration, often considered a homogeneous coalition, leads to stable and cost-effective matching policies. Furthermore, the non-homogeneous formations can also be a feasible and comparatively fair candidate solution for both sub-figures. In fact, most of the coalitions will fit into the core game of E-CPDPTW and deploy at least one UAV in the fleet; the requests are delivered in such a network size, yet not as efficient as operating more ADRs to deal with the high-capacity demands. In other words, in delivery systems, especially in large networks, it is more convenient if one UAV goes to multiple pickup and delivery locations than if one of them is assigned to an ADR if not mandated by capacity constraint violation. Moreover, the battery consumption rate of UAVs is significantly higher than that of ADRs compared to the current state-of-the-art technology, which leads them to visit recharging stations more frequently and puts more load on urban areas’ serviceability. Therefore, using coalitions of ground mode, which can keep away from this shortcoming, is of paramount importance, as can be implied from the result of the core analysis.
@@
Finally, according to Figure 16, the optimal coalition for both cases is not when all resources are available; instead, middle regions of the plot, where starting from 3-5 UAVs and 3-5 ADRs, tend to exhibit higher gain. Although it has been inferred that having more vehicles will substantiate the core existence and applicability, it can be achieved in the case of $n=120$ with only $80\%$ of resources. This is also observable in the case of $n=50$ when ADRs are of the $80\%$ of resources. Certain scenarios as well such as $r=4,d=\{2,3\}$ on the scale of 50 or $r=3,d=\{1,2,3\}$ on the scale of 120, do not support the overall trend of the graph, which tend to gain more values for utilizing more vehicles. The rationale behind this is mainly due to a misfit in the generalization of larger to lower fleets, which causes the graph not to be smooth and continuously changing. However, the explanation for the former statement of maximum utilization gain is believed to be underlying the fact that systems have found that specific coalition to be more optimal given the prioritizing of the sufficient number of UAVs to avoid the high cost. Note that not necessarily adding more vehicles will lead to redundancy, where each additional agent contributes less benefit to the coalition and may put extra effort into other vehicles and reduce the overall efficiency. As a result, achieving the optimal coalition value depends on finding the right balance between the number of UAVs and ADRs.
@@ 6 Conclusions and Future Directions
We present a comprehensive framework for multi-modal autonomous delivery utilizing ADRs and UAVs in urban environments. By addressing critical challenges related to on-demand delivery in high-density areas and efficient vehicle allocation, the proposed framework aims to enhance the accessibility of delivery services while optimizing network infrastructure. Developing an interactive cooperative allocation system ensures a centralized decision-making process to determine the most effective delivery mode regarding efficiency and time delay.
@@
The stochastic nature of the delivery problem is acknowledged, considering factors like unknown demand, airspace regulations, and unpredictable weather conditions. Our solution tailored a deep reinforcement learning approach, specifically the transformer architecture with edge-enhanced attention models, to achieve optimal assignment and routing in urban environments. By simulating the delivery task for the fleet, our methodology accounts for delivery priority, vehicle states, and environmental changes. Moreover, the coalitional game theory model has been adopted to investigate potential cooperative cost profit by which various coalitions are formed. In addition, core properties where they proved to exist have been established to identify multi-modal delivery advantages over uni-modal and cost distribution among groups of vehicles.
@@
Our proposed model exhibits superior performance in urban on-demand delivery systems applied in various delivery scenarios, giving importance from customers’ to commercial perspective to accommodate large high-density areas utilizing cooperative operation of UAVs where reducing delay costs and handling time-sensitive deliveries is accomplished and ADRs carrying larger payloads and decrease the burden from the fleet is done. The problem of E-CPDPTW has been investigated in a variety of aspects, including robustness tests of operating uncertain and irregular patterns of time-location distribution as well as wind conditions, which could be potentially utilized in such situations owing to its sustainable performance. In addition, coalition following the generalization ability of such a mode has been analyzed in various configurations, scales, and multiple agents, verifying a robust and scalable model.
@@
There are several future directions where the research can be further developed. Firstly, refinement of the proposed framework will be essential to account for real-time constraints and dynamic adjustments in response to changing environmental conditions, such as congestion in the edge insertion technique. Exploration of additional factors, such as airspace conflict resolution and local regulations, will contribute to a more realistic and adaptable delivery system. Additionally, the system’s scalability in the context of multi-agent reinforcement learning warrants investigation. Extending the framework to handle a larger fleet of autonomous vehicles, diverse urban environments of different layouts, and decentralized execution will be crucial for practical implementation and widespread adoption. Furthermore, evaluating the proposed on-demand delivery system under varying external factors, such as vehicle failure and order cancellation, will be a focus of future work. Continuous optimization of the objective function, considering both the level of service and delivery cost, will contribute to an increasingly efficient and resilient autonomous delivery system.
Appendix A Mixed Integer Programming (MIP) Model for E-CPDPTW
The objectives and constraints in the mathematical formulation are listed as follows. The objective function is shown in the Equation 30.
$\text{min}\;R=\alpha_{1}\sum_{k\in N^{d}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\alpha_{2}\sum_{k\in N^{r}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\sum_{i\in P\cup D}\alpha_{3}\left|T_{ik}-e_{i}\right|+\sum_{i\in P\cup D}max\left\{T_{ik}-l_{i},0\right\}$ (30)
Where $t_{ijk}$ is time travelled by vehicle $k$ between node $i$ and $j$, $T_{ik}$ is the arrival time of the vehicle $k$ at node $i$, $\alpha_{1}$ and $\alpha_{2}$ are positive monetary conversion factors pf utilizing UAVs and ADRs respectively per hour. Also, the $\alpha_{3}$ is the monetary coefficient of waiting time penalty for customers that the delivery company has to compensate in the case of delayed delivery. The second and third terms can only be positive because of the violation of time constraints. If, in the second term, the vehicle happens to arrive earlier than the specified time window, we treat the time difference loss as the first term of the cost function. Additionally, the equations that constrain this objective function are discussed below.
$\sum_{j=0}^{P\cup D}X_{ijk}=1,\forall k\in\left\{N^{r}\cup N^{d}\right\}\;\forall i\in C$ (31)
$\sum_{i=0}^{P\cup D}X_{ijk}=1,\forall k\in\left\{N^{r}\cup N^{d}\right\}\;\forall j\in C$ (32)
$\sum_{k=1}^{N^{r}\cup N^{d}}\sum_{i=0}^{N}X_{ijk}=1,\forall i\in\{P\cup D\}$ (33)
$\sum_{i=0}^{P\cup D}X_{ijk}=\sum_{i=0}^{P\cup D}X_{i\left(j+n\right)k},\forall j\in P,\;\forall k\in\left\{N^{r}\cup N^{d}\right\}$ (34)
$T_{ik}\geq e_{i},\forall k\in\left\{N^{r}\cup N^{d}\right\},\forall i\in\{P\cup D\}$ (35)
$T_{ik}\leq T_{k,i+n},\forall i\in\left\{P\right\},\forall k\in\left\{N^{r}\cup N^{d}\right\}$ (36)
$T_{ki}+s_{i}+t_{ijk}\leq T_{kj}+(1-M)X_{ijk},\forall k\in\left\{N^{r}\cup N^{d}\right\},\forall(i,j)\in\{P\cup D\}$ (37)
$0\leq u^{k}_{i}\leq Q_{k},\forall k\in\left\{N^{r}\cup N^{d}\right\},\forall i\in\{P\cup D\}$ (38)
$u^{k}_{i}=u^{k}_{j}+q_{j}+(1-M)X_{ijk},\forall k\in\left\{N^{r}\cup N^{d}\right\},\forall(i,j)\in\{P\cup D\}$ (39)
$e^{k}_{i}-e_{ijk}+B^{k}\left(1-X_{ijk}\right)\geq e_{j},\forall k\in\left\{N^{r}\cup N^{d}\right\},\forall(i,j)\in\{P\cup D\}$ (40)
$e_{min}\leq e^{k}_{i}\leq B^{k},k\in\left\{N^{r}\cup N^{d}\right\},\forall i\in\{P\cup D\}$ (41)
$e^{k}_{i}=B^{k},\forall k\in\left\{N^{r}\cup N^{d}\right\},i\in C$ (42)
$X_{ijk}\in\{0,1\},\forall(i,j)\in\{P\cup D\},\forall k\in\left\{N^{r}\cup N^{d}\right\}$ (43)
The constraints in Equations 31 and 32 ensure the vehicles depart from and return to any depot, whether for recharging or ending the tour. The Equations 33 and 34 ensure the requests are only served once and by the same vehicle. The constraints in Equations 35, 36, 37 ensure that vehicles cannot start service at a location earlier than its early time window and cannot be later than the service at its corresponding delivery location, and $M$ is a positive large number. All vehicles must service the pickup location earlier than the delivery node and account for each node’s service time, $s_{i}$. The capacity constraint of the vehicles and the load balance update is given in Equations 38 and 39, with $u^{k}_{i}$ denoting the $k$ th vehicle’s load after the service at the $i$th location is done. Equations 40, 41, and 42 denote battery level balance update, capacity constraint, and recharging state after visiting depots, respectively, with $e^{k}_{i}$ denoting the $k$ th vehicle’s battery level after going to $i$ th location. Lastly, Equation 43 notes that the binary variable.
Table 5: Parameters and Definitions
Parameter Definition
$C$ Depot instances set
$P$ Set of pickup nodes
$D$ Set of Delivery nodes
$E^{d}$ Set of UAV’s edges network
$E^{r}$ Set of ADR’s edges network
$[e_{i},l_{i}]$ Early and late time window
$q_{i}$ Demand of customer
$N^{d}$ Number of UAVs
$N^{r}$ Number of ADRs
$N^{d,r}$ Number of vehicles
$Q_{k}$ Capacity of vehicle
$B_{k}$ Battery level of vehicle
$N$ Number of customers
$T_{ik}$ Arrival time of the vehicle $k$ at node $i$
$t_{ijk}$ time traveled by vehicle $k$ between node $i$ and $j$
$e_{min}$ The minimum battery capacity of vehicle $j$
$X_{ijk}$ Binary decision variable for whether the vehicle $k$ travels from vertex $i$ to vertex $j$
Appendix B UAV Energy Consumption
@@
where n is the number of rotors, and $\varsigma$ is the area of the spinning blade disc of one rotor. Besides, the angle of attack $\alpha$ and thrust are given by Equation 46 and 47, respectively.
@@
Appendix C Training Algorithm

### v2

@@ 5.5.1 Coalition Generalizability
The overall result shows an acceptable solution generated from each of the scales to others; most notably, the generalization on smaller networks from the trained model of larger networks produces better results. This is probably due to the inclusion of smaller and more information on the network embedding. This variation can also be explained by the different number of depots in each case, which causes the training parameter not to capture the depots’ embedding well. In addition, as the scale of the problem gets larger, generalization results undergo variability, particularly due to managing the fewer vehicles within the larger network and generalizing from lower vehicle-to-request ratio distributions, which causes ineffectiveness in high request densities assignment with limited resources. Nevertheless, in the constant size of the graph, the model’s ability to generalize is still effective, and in the next part, we analyzed the core game theory for the constant size of the network, with various numbers of agents.
@@ 5.5.2 Coalitional Ability
We use the pre-trained model for E-CPDPTW120-d3 for ten agents (five of each mode) to conduct a test to find the core coalition among the subset of agents. To evaluate the cost function for each coalition, we use generalization on the pre-trained model to obtain cost values to avoid the high computational expense of training for each coalition. The coalitions are of the two separate modes of ADRs ($r\in\left\{1,\ldots,N^{r}\right\}$) and UAVs ($d\in\left\{1,\ldots,N^{d}\right\}$) set, assumed to have no intersection.
@@
The Algorithm 1 is used here to assess how coalitions will perform and if the core exists. For a group consisting of either UAVs or ADRs, the individual cost of both modes will be calculated, as well as their coalition. Subsequently, the sub-additivity and efficient conditions are checked so that not only must the cooperative cost be less than both individual modes, but also, the newly allocated cost of the coalition must be less than the average mode cost alone. Figure 15 displays contour plots for coalition gains; the heat map bar shows the difference between the sum of individual and cooperative modes. Note that the spectrum level is set to be sufficiently high to obtain a smooth colour transition graph. However, it is clear that the core analysis plot is discrete. The star sign in the figure indicates if the core exists, given its criteria. It was observed in both scales of the problem that the core exists except in the network size of 120, where only one UAV is operating. This can be mainly because in the larger network, having one UAV independently from the ADR numbers cannot accommodate the demand due to high delays in the deliveries of ADRs, yet it is achievable in the $n=50$. Nonetheless, lower coalition values are observed in both graphs where fewer UAVs are involved, and the lowest value gain is obtained, which fails to perform well in such a large network. It is, however, toward yellowish regions where the most value is gained by acting cooperatively, which is increasing where both the number of UAVs and ADRs increase. This implies that a balanced configuration, often considered a homogeneous coalition, leads to stable and cost-effective matching policies. Furthermore, the non-homogeneous formations can also be a feasible and comparatively fair candidate solution for both sub-figures. In fact, most of the coalitions will fit into the core game of CE-CPDPTW and deploy at least one UAV in the fleet; the requests are delivered in such a network size, yet not as efficiently as operating more ADRs to deal with the high-capacity demands. In other words, in delivery systems, especially in large networks, it is more convenient if one UAV goes to multiple pickup and delivery locations than if one of them is assigned to an ADR if not mandated by a capacity constraint violation. Moreover, the battery consumption rate of UAVs is significantly higher than that of ADRs compared to the current state-of-the-art technology, which leads them to visit recharging stations more frequently and puts more load on urban areas’ serviceability. Therefore, using coalitions of ground mode, which can keep away from this shortcoming, is of paramount importance, as can be implied from the result of the core analysis.
Additionally, the cost allocation by Shapley value is illustrated in Figure 16, distribution of share for each mode across each coalition is shown. The cost share over all coalition is fixed since the UAV and ADR fleet are homogeneous and they represent a higher gain for UAVs on dominance conditions since they provide higher marginal value whereas ADRs are less sensitive to scale and are more efficient in terms of cost sharing in smaller networks. Next, a case of non-homogeneous demand distribution is considered, where payload sizes are from $(3,8)$ bound as well as a gaussian location distribution to experiment the case of deliveries are clustered in dense area. The additivity check and cost allocation are shown in Figure 17. In this Figure, the potential marginal contribution for UAVs is subjected to more reducing share in cost due to more load on ADRs. Specifically, they are more effective in coalitions which in case of uniform data distribution was more UAV-oriented. Additionally, cost shares in ADR shows faster as number of agent which accounts for marginal contribution of ADR in such scneraios.
(a) Coalitional analysis for the network size of 50.
(b) Coalitional analysis for the network size of 120.
@@
(a) Shapely value for network size of 50.
(b) Shapely value for the network size of 120.
Figure 16: Cost allocation for all coalitions. The first number on the left show UAV and second ADR.
(a) Coalitional analysis for the network size of 50.
(b) Shapely value for the network size of 50.
Figure 17: Cost allocation for non-homogeneous demand distribution.
Finally, according to Figure 15, the optimal coalition for both cases is not when all resources are available; instead, the middle regions of the plot, where starting from 3-5 UAVs and 3-5 ADRs, tend to exhibit higher gain. Although it has been inferred that having more vehicles will substantiate the core existence and applicability, it can be achieved in the case of $n=120$ with only $80\%$ of resources. This is also observable in the case of $n=50$ when ADRs are of $80\%$ of resources. Certain scenarios, such as $r=4,d=\{2,3\}$ on the scale of 50 or $r=3,d=\{1,2,3\}$ on the scale of 120, do not support the overall trend of the graph, which tends to gain more values for utilizing more vehicles. The rationale behind this is mainly due to a misfit in the generalization of larger to smaller fleets, which causes the graph not to be smooth and continuously changing. However, the explanation for the former statement of maximum utilization gain is believed to be underlying the fact that systems have found that specific coalition to be more optimal, given the prioritizing of a sufficient number of UAVs to avoid the high cost. Note that not necessarily adding more vehicles will lead to redundancy, where each additional agent contributes less benefit to the coalition and may put extra effort into other vehicles and reduce the overall efficiency. As a result, achieving the optimal coalition value depends on finding the right balance between the number of UAVs and ADRs. This balance can be affected by different test cases, and each mode share can vary significantly, as the recent analysis shows it will be more ADR or UAV needed for larger scale, and in general, the more agents involved, the more efficient the cost allocation results, in addition to the cases which both can complement their delivery task which the other cannot achieve.
@@ 6 Conclusions and Future Directions
We present a comprehensive framework for multi-modal autonomous delivery utilizing ADRs and UAVs in urban environments. By addressing critical challenges related to on-demand delivery in high-density areas and efficient vehicle allocation, the proposed framework aims to enhance the accessibility of delivery services. Developing an interactive cooperative allocation system ensures a centralized decision-making process to determine the most effective delivery mode regarding efficiency and time delay. The stochastic nature of the delivery problem is acknowledged, considering factors such as a heterogeneous fleet interaction in urban areas cluttered with buildings, airspace regulations, and unpredictable weather conditions. Our solution tailored a deep reinforcement learning approach, specifically the transformer architecture with edge-enhanced agent-aware attention models capturing spatial and temporal coupling effects, to achieve optimal assignment and routing in urban environments. By simulating the delivery task for the fleet, our methodology accounts for delivery priority, vehicle states, and environmental changes. Moreover, the coalitional game theory model has been adopted to investigate potential cooperative cost-profit by which various coalitions are formed. In addition, core properties where they proved to exist have been established to identify multi-modal delivery advantages over uni-modal and cost distribution among groups of vehicles.
@@
Our proposed model exhibits superior performance in urban on-demand delivery systems applied in various delivery scenarios, giving importance from customers’ to commercial perspectives, utilizing potential cooperative operation of UAVs, where reducing delay costs and handling time-sensitive deliveries is accomplished, and ADRs carrying larger payloads and decreasing the burden on the fleet is done. The problem of CE-CPDPTW has been investigated in a variety of aspects, including robustness tests of operating uncertain and heterogeneous patterns of time-location distribution as well as wind conditions, which could be potentially utilized in such situations owing to its sustainable performance. In addition, the coalition following the generalization ability of such a mode has been analyzed in various configurations, scales, and multiple agents, demonstrating the cooperative benefit in large-scale and non-homogeneous demand distribution.
@@
There are several future directions where the research can be further developed. Firstly, the centralized conflict resolution three-dimensional network plan to manage the UAV traffic throughout the vertical and horizontal aerial links to avoid potential collisions. Exploration of additional factors, such as non-homogeneous sub-fleet vehicles ranging in different capacities and battery sizes, will contribute to a more realistic and prioritized order in the delivery system. Additionally, the system’s scalability in the context of multi-agent systems by extending the framework to handle a larger fleet of autonomous vehicles, diverse urban environments of different layouts, and decentralized execution will be crucial for practical implementation and widespread adoption. Furthermore, evaluating the proposed on-demand delivery system under varying external factors, such as vehicle failure and split delivery, by incorporating the adaptive generalization techniques, will be a focus of future work.
Appendix A UAV Energy Consumption
@@
where n is the number of rotors, and $\varsigma$ is the area of the spinning blade disc of one rotor. Besides, the angle of attack $\alpha$ and thrust are given by Equation 53 and 54, respectively.
@@
Appendix B Training Algorithm